ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60,90 Nummelin, Esa, University of Helsinki, Finland Large deviations of random vector fields with applications to economics We consider the large deviations of price equilibria for random economic systems. The large deviation rates are characterized with the aid of an entropy function I(p) where the price vectors p act as “intensive” thermodynamic variables. Reference: E. Nummelin, On the existence and convergence of price equilibria for random economies. Manuscript (1998), submitted for publication.